OpenStudy (anonymous):

Can anyone help me to use the quadratic formula to solve this equation? 4x^2 -9x -1 = 0

OpenStudy (jdoe0001):

the quadratic formula is pretty straight forward just plug your values in

OpenStudy (anonymous):

I have and I have an answer, I just want to check if my answer is correct.

OpenStudy (anonymous):

i will check your answer what is it :)

OpenStudy (anonymous):

x=9 + sqrt. 97/8, and 9 - sqrt 97/8

OpenStudy (nurali):

4x^2 -9x -1 = 0 a=4, b=-9 and c=-1 x=-b+-sqrt(b^2-4ac)/2a x=9+-sqrt(81+16)/8 x=9+-sqrt(97)/8 x=9 + sqrt. 97/8, and 9 - sqrt 97/8

OpenStudy (anonymous):

Nurali you got the same answer as I did so this is correct right? because I also thought it might be x = 1/8 (9-sqrt 97) x = 1/8 (9+ sqrt 97)

OpenStudy (nurali):

Both are right.

OpenStudy (anonymous):

Cool! You are racking up medals!